A firm has debt of $5,000, equity of $16,000, a cost of debt of 8 percent, a cost of equity of 12 percent, and a tax rate of 21 percent. What is the firm's weighted average cost of capital?

10.20 percent

9.94 percent

10.90 percent

10.65 percent

11.05 percent
Answer #1

After-tax cost of debt=8*(1-tax rate)

=8*(1-0.21)=6.32%

Total=(5000+16000)=$21000

WACC=Respective cost*Respective weight

=(5000/21000*6.32)+(16000/21000*12)

=10.65%(Approx)